Calculate the lattice energy for the formation of MgO from Mg and O 2
given the following information:
½ O 2 = O 249 kJ
O + 1 e - = O - -141 kJ
O -1 + 1 e - = O -2 878 kJ
Mg = Mg +1 + 1 e - 738 kJ
Mg +1 = Mg +2 + 1 e - 1450 kJ

Mg +2 + O -2 = MgO -601 kJ
1
Expert's answer
2019-07-25T05:29:51-0400

Let's sum up

1/2O2+O+1e+O-+1e+Mg+Mg1++Mg2++O2-=O+1e+O-+1e+Mg+Mg1++Mg2++O2--1803kJ

1/2 O2+Mg=MgO -1803kJ

∆H(MgO) =1803 kJ